Prove that n=2 mod4,and when n=p ,where p is a prime and p≡3 mod4

1
Expert's answer
2021-12-13T14:00:18-0500

"n \equiv 3(\bmod 4)," therefore all the prime factors of n are odd.

For contradiction, assume all of the prime factors are equivalent to "1 \bmod 4."

Then prove n is also equivalent to "1 \bmod 4" , which is a contradiction (note "(4 k+1)(4 m+1)=4(k m+k+m)+1)."

 Suppose for contradiction there are only finitely many primes of the form 4k+3, and call the set of all of them "\left\{p_{1}, p_{2}, \ldots, p_{n}\right\}." But then

"\left(p_{1} p_{2} \cdots p_{n}\right)^{2}+2=3 \quad(\bmod 4)"

by (1) has a prime factor "p_{n+1}" equivalent to "3 \bmod 4" .

By definition of "\left\{p_{1}, p_{2}, \ldots, p_{n}\right\}" , we get "p_{n+1} \in\left\{p_{1}, p_{2}, \ldots, p_{n}\right\},"

Thus "p_{n+1}\left|\left(p_{1} p_{2} \cdots p_{n}\right)^{2}+2 \Longrightarrow p_{n+1}\right| 2"

Contradiction.